OpenStudy (anonymous):

How can I solve e^(2x)=1/e^(x+6)?

OpenStudy (anonymous):

Can I make the right side of the equation into e^(-x+6) to get rid of the fraction?

OpenStudy (raffle_snaffle):

Do you remember your ln and log rules?

OpenStudy (raffle_snaffle):

example  to eliminate e you need to take the ln of both sides

OpenStudy (raffle_snaffle):

Answer to your question you can

OpenStudy (campbell_st):

start by rewriting the right hand side in index form \[e^{2x} = e^{-(x + 6)}\] now you have the same base, you can equate the powers and solve for x hope this helps

OpenStudy (anonymous):

So would I have 2x=x+6 or 2x=-x-6?

OpenStudy (anonymous):